40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 150 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  9533   Tue Jan 7 23:13:47 2014 jamieUpdateCDS/frames is full, causing daqd to die

Quote:

So why is /frames full?  Apparently the wiper script is either not running, or is failing to do it's job.  My guess is that this is a side effect of the linux1 raid failure we had over xmas.

It actually looks like the wiper script has been running fine.  There is a log from Tuesday morning:

controls@fb ~ 0$ cat /opt/rtcds/caltech/c1/target/fb/wiper.log

Tue Jan  7 06:00:02 PST 2014

Directory disk usage:
/frames/trend/minute_raw 385289132k
/frames/trend/second 100891124k
/frames/full 12269554048k
/frames/trend/minute 1906772k
Combined 12757641076k or 12458633m or 12166Gb

/frames size 13460088620k at 94.78%
/frames is below keep value of 95.00%
Will not delete any files
df reported usage 97.72%
controls@fb ~ 0$

So now I'm wondering if something else has been filling up the frames today.  Has anything changed today that might cause more data than usual to be written to frames?

I'm manually running the wiper script now to clear up some /frames.  Hopefully that will solve the problem temporarily.

  9532   Tue Jan 7 23:09:10 2014 manasaUpdateIOOMC aligned

Quote:

[Rana, Jenne]

We turned off the WFS servos, and looked at the MC REFL DC, and saw that it was still good, so we said that since the MC spots are pretty good, that we'll keep this alignment for now. 

Rana put the beam back on the center of the IOO QPDs on the PSL table.

We switched a steering mirror in the WFS path that was the wrong handed-ness to be the correct handed-ness, then put the beam on the centers of the WFS.  We turned on the WFS, and everything seems good. 

There were no major drifts in the WFS error signals while we were gone for dinner, so the MC seems okay for now.

 The last 4 hour trend for WFS error signals show some amount of drift. We should still look at the long term trend to solve the issue.

Attachment 1: WFSdrift.png
WFSdrift.png
  9531   Tue Jan 7 23:08:01 2014 jamieUpdateCDS/frames is full, causing daqd to die

Quote:

The daqd process is segfaulting and restarting itself every 30 seconds or so.  It's pretty frustrating. 

Just for kicks, I tried an mxstream restart, clearing the testpoints, and restarting the daqd process, but none of things changed anything.  

Manasa found an elog from a year ago (elog 7105 and preceding), but I'm not sure that it's a similar / related problem.  Jamie, please help us

The problem is not exactly the same as what's described in 7105, but the symptoms are so similar I assumed they must have a similar source.

And sure enough, /frames is completely full:

controls@fb /opt/rtcds/caltech/c1/target/fb 0$ df -h /frames/
Filesystem            Size  Used Avail Use% Mounted on
/dev/sda1              13T   13T     0 100% /frames
controls@fb /opt/rtcds/caltech/c1/target/fb 0$

So the problem in both cases was that it couldn't write out the frames.  Unfortunately daqd is apparently too stupid to give us a reasonable error message about what's going on.

So why is /frames full?  Apparently the wiper script is either not running, or is failing to do it's job.  My guess is that this is a side effect of the linux1 raid failure we had over xmas.

  9530   Tue Jan 7 22:44:45 2014 JenneUpdateCDSdaqd on fb is segfaulting every ~30 seconds

The daqd process is segfaulting and restarting itself every 30 seconds or so.  It's pretty frustrating. 

Just for kicks, I tried an mxstream restart, clearing the testpoints, and restarting the daqd process, but none of things changed anything.  

Manasa found an elog from a year ago (elog 7105 and preceding), but I'm not sure that it's a similar / related problem.  Jamie, please help us!

Here is a screen dump from the "dtail":

Every 1.0s: dmesg | tail -50                                                                                                                         Tue Jan  7 22:43:23 2014

[   33.498691]  [<ffffffff8104a063>] kthread+0x7a/0x82
[   33.498695]  [<ffffffff81003654>] kernel_thread_helper+0x4/0x10
[   33.498698]  [<ffffffff81049fe9>] ? kthread+0x0/0x82
[   33.498701]  [<ffffffff81003650>] ? kernel_thread_helper+0x0/0x10
[   33.498703] ---[ end trace 6236defa99b3e091 ]---
[   33.498705] mx INFO: Board 0: allocated MSI IRQ 67
[   33.498713] mx INFO: CPU0: PAT = 0x7010600070106
[   33.498715] mx INFO: CPU0: new PAT = 0x1010600070106
[   33.498718] mx INFO: Board 0: Using PAT index 6
[   33.499101] eth0: no IPv6 routers present
[   33.531013] mx INFO: Board 0: device 8, rev 0, 1 ports and 2096896 bytes of SRAM available
[   33.531017] mx INFO: Board 0: Bridge is 10de:005d
[   33.531228] mx INFO: Board 0: MAC address = 00:60:dd:46:ea:ec
[   33.535971] mx INFO: Loaded mcp of len 235448
[   34.489244] mx INFO: Starting usermode mapper at /opt/mx/sbin/mx_start_mapper
[   39.148855] mx INFO: mx0: Link0 is UP
[   39.588511] mx INFO: myri0: Will use skbuf frags (4096 bytes, order=0)
[   39.589299] mx INFO: 1 Myrinet board found and initialized
[  287.706367] daqd used greatest stack depth: 3368 bytes left
[86605.907520] daqd[18407]: segfault at 38b08e4c0 ip 00007f11b3942a6c sp 00007f10b1917d50 error 4
[86605.907530] daqd[18424]: segfault at 38b544f90 ip 00007f11b3942a6c sp 00007f10b12c6d30 error 4 in libc-2.10.1.so[7f11b390e000+14c000] in libc-2.10.1.so[7f11b390e000+14c00
0]
[86605.907544]
[86605.919454] daqd[21319] general protection ip:7f11b3942a6c sp:7f10b1814d30 error:0
[86605.919462] daqd[18442] general protection ip:7f11b3942a6c sp:7f10b0bf4d30 error:0
[86605.919615] daqd[18443]: segfault at 38aee3db0 ip 00007f11b3942a6c sp 00007f10b0b73d50 error 4 in libc-2.10.1.so[7f11b390e000+14c000]
[86605.919694] daqd[18412]: segfault at 38aff35d0 ip 00007f11b3942a6c sp 00007f10b1752d30 error 4
[86605.919701] daqd[18417]: segfault at 38b544f70 ip 00007f11b3942a6c sp 00007f10b154dd50 error 4 in libc-2.10.1.so[7f11b390e000+14c000]
[86605.919708] daqd[18445]: segfault at 38aff35b0 ip 00007f11b3942a6c sp 00007f10b0ab1d50 error 4
[86605.919733] daqd[18429]: segfault at 38b42ae90 ip 00007f11b3942a6c sp 00007f10b10c1d50 error 4 in libc-2.10.1.so[7f11b390e000+14c000]
[86605.919741] daqd[18440]: segfault at 38b08e480 ip 00007f11b3942a6c sp 00007f10b0cb6d30 error 4 in libc-2.10.1.so[7f11b390e000+14c000]
[86605.958551]  in libc-2.10.1.so[7f11b390e000+14c000] in libc-2.10.1.so[7f11b390e000+14c000]
[86605.958557]
[86605.958577]  in libc-2.10.1.so[7f11b390e000+14c000]
[86605.958586]  in libc-2.10.1.so[7f11b390e000+14c000]
[86605.959639] daqd used greatest stack depth: 3160 bytes left
[98139.100888] show_signal_msg: 13 callbacks suppressed
[98139.100895] daqd[23753]: segfault at 39c7363b0 ip 00007f5bf253ba6c sp 00007f5b69b48d30 error 4 in libc-2.10.1.so[7f5bf2507000+14c000]
[98687.815120] daqd used greatest stack depth: 2984 bytes left
[208995.594227] daqd[10386] general protection ip:7f3b7c930a6c sp:7f3a79f09d50 error:0 in libc-2.10.1.so[7f3b7c8fc000+14c000]
[353015.067479] daqd used greatest stack depth: 2880 bytes left
[367406.863618] daqd[13078]: segfault at 41 ip 0000000000000041 sp 00007fb1f0ba2cf8 error 14 in daqd[400000+7c000]
[367406.863833] daqd[13104] general protection ip:7fb2f3018a6c sp:7fb1f01c8d30 error:0
[367406.863877] daqd[13086] general protection ip:7fb2f3018a6c sp:7fb1f089ad30 error:0
[367406.877408] daqd[13080]: segfault at 41 ip 0000000000000041 sp 00007fb1f0ae0ca8 error 14 in daqd[400000+7c000]
[367406.877435]  in libc-2.10.1.so[7fb2f2fe4000+14c000]
[367406.877442] daqd[13100]: segfault at 39ba287b0 ip 00007fb2f3018a6c sp 00007fb1f034cd30 error 4 in libc-2.10.1.so[7fb2f2fe4000+14c000]
[367406.878372]  in libc-2.10.1.so[7fb2f2fe4000+14c000]
[399802.887523] daqd[18295] general protection ip:7fb056a71a6c sp:7faf96125f10 error:0 in libc-2.10.1.so[7fb056a3d000+14c000]
[410595.969327] daqd[22057]: segfault at 3a91f27b0 ip 00007f48e96eea6c sp 00007f47e6c26d50 error 4 in libc-2.10.1.so[7f48e96ba000+14c000]
[410595.988926] daqd[22068]: segfault at 3a91f2790 ip 00007f48e96eea6c sp 00007f47e681bd30 error 4 in libc-2.10.1.so[7f48e96ba000+14c000]

  9529   Tue Jan 7 21:00:02 2014 JenneUpdateIOOIP POS, IP ANG aligned

After locking the arms (after the MC alignment work), Manasa and I aligned IP POS, IP ANG, and both end transmission QPDs. 

We noticed that IP ANG is clipping in yaw as it comes onto the end table.  It looks to me like it's clipping on the edge of the plastic box's aperture, but I can't guarantee that it's not also clipping elsewhere. 

  9528   Tue Jan 7 20:57:41 2014 JenneUpdateIOOMC aligned

[Rana, Jenne]

We turned off the WFS servos, and looked at the MC REFL DC, and saw that it was still good, so we said that since the MC spots are pretty good, that we'll keep this alignment for now. 

Rana put the beam back on the center of the IOO QPDs on the PSL table.

We switched a steering mirror in the WFS path that was the wrong handed-ness to be the correct handed-ness, then put the beam on the centers of the WFS.  We turned on the WFS, and everything seems good.

While we were out on the table, we also changed the anodized aluminum dump for a razor dump, to catch the reflection from the 2inch lens that is the first thing the MC refl path sees out of vac.

There were no major drifts in the WFS error signals while we were gone for dinner, so the MC seems okay for now.

  9527   Tue Jan 7 17:16:04 2014 manasaUpdateIOOMC aligned

Quote:

Edit, JCD:  What this really means is that our DC MC pointing is bad, and we need to move the MC suspensions to offload the WFS.  (All of the WFS output numbers for MC1 and 3 were around 100, which is pretty big for those numbers).  We should resurrect the WFS offloading scripts so that we can do this more regularly, and not have to do it by hand.

 Aligned MC to offload the WFS

1. Turned OFF the WFS feedback servo.

2. Aligned the MC suspensions by moving the pit and yaw sliders. MC trans sum brought from ~11000 counts to ~15000 counts. MC RFPD DCMON reads 0.45 counts.

3. Turned ON the WFS servo. The WFS output now reads in the order of 0 to +/-15.

4. Measured the MC spot positions. The spot positions look like they moved for the better compared to what they were yesterday.

 

Attachment 1: MCspots.png
MCspots.png
  9526   Tue Jan 7 16:41:08 2014 manasaUpdateIOOWFS moving MC suspensions

Quote:

 The trend shows a big jolt to the MC1/3 pointing this morning at 8:30.

Was anyone working anywhere near there today? There is no elog.

If not, we will have to put a 'no janitor' sign on all of the 40m doors permanently to prevent mops misaligning our interferometer.

The MC trend for the last 2 days shows that the MC suspensions were kicked again earlier today.  Looking back at the suspension channel INMONs along with the MC trans sum shows that the suspensions get kicked everytime MC locks and unlocks. (Attch:1)

So I checked the effect of WFS on the suspensions by disabling and enabling the WFS feedback servo (Attch:2).

Since the IMC is not at it best pointing, whenever the  MC autolocker runs and enables the WFS, the suspensions look like they are getting kicked.  But really, it's just the WFS doing their job. 

Edit, JCD:  What this really means is that our DC MC pointing is bad, and we need to move the MC suspensions to offload the WFS.  (All of the WFS output numbers for MC1 and 3 were around 100, which is pretty big for those numbers).  We should resurrect the WFS offloading scripts so that we can do this more regularly, and not have to do it by hand.

Attachment 1: 2dayMCtrend.png
2dayMCtrend.png
Attachment 2: WFSvsMCsuspensions.png
WFSvsMCsuspensions.png
  9525   Tue Jan 7 11:11:36 2014 ranaUpdateIOOMC drift

 

 NOT drift. The sudden steps are certainly the result of being kicked. The slow drift at the end of the day might be a slow strain relaxation.

It pays to be careful and not put too much weight or impulsive forces on the chambers or tables.

  9524   Tue Jan 7 10:44:13 2014 SteveUpdateIOOMC drift

Quote:

 The trend shows a big jolt to the MC1/3 pointing this morning at 8:30.

Was anyone working anywhere near there today? There is no elog.

If not, we will have to put a 'no janitor' sign on all of the 40m doors permanently to prevent mops misaligning our interferometer.

 I was taking pictures at the AP table at the morning and ETMX optical table after noon. There was no activity on the IOO chamber.

 Look at the last 2 hours of  Rana's trend plot. MC1 and MC2 sensor voltage started increasing.

I think it was a drift action.

Attachment 1: 2dTrend.png
2dTrend.png
Attachment 2: driftNotKick.png
driftNotKick.png
  9523   Mon Jan 6 22:11:46 2014 JenneUpdateLSCPRCL sideband locking still not so happy

 

 The PRCL once again doesn't want to lock on sidebands for me.  I can lock on the carrier just fine (using the IFO Config settings, along with some hand-alignment of the PRM). 

However, I can't convince it to lock on sidebands.  Using the configs that I used on Dec 18th (elog 9491), I'm not getting it.  I've done the arm ASS alignment, and I've run LSCoffsets, both of which seemed to do their things appropriately. 

I'm going to attribute this today to not being in the groove yet, and I'll look at it again in the morning.

  9522   Mon Jan 6 20:52:09 2014 JenneUpdateIOOMC1/3 kicked this morning at 8:30

When I got in this morning at 9-something (9:45 maybe?), Steve was taking dust photos on the AS table, of the MC Refl path.  Other than that, I don't have any information. 

Also, Tuesday is our traditional janitor day, so I'm hesitant to put our blame there.  (I think we've kept Tuesdays, even though we're on a less-often schedule....Steve will have to correct me if I'm wrong on this).

  9521   Mon Jan 6 18:32:17 2014 RANAUpdateIOOMC1/3 kicked this morning at 8:30

 The trend shows a big jolt to the MC1/3 pointing this morning at 8:30.

Was anyone working anywhere near there today? There is no elog.

If not, we will have to put a 'no janitor' sign on all of the 40m doors permanently to prevent mops misaligning our interferometer.

Attachment 1: kicked.png
kicked.png
  9520   Mon Jan 6 16:32:40 2014 KojiSummaryGenerallinux1 RAID crash & recovery

Since this configuration change, the daily backup was speeded up by factor of more than two.
It was really limited by the bandwidth of the RAID array.

/cvs/cds/rtcds/caltech/c1/scripts/backup/rsync.backup.cumlog:

...
rsync.backup start: 2013-12-20-05:00:00, end: 2013-12-20-07:04:28, errcode 0
...
rsync.backup start: 2014-01-05-05:00:00, end: 2014-01-05-05:55:04, errcode 0

(The daily backup starts from 5:00)

  9519   Mon Jan 6 16:30:31 2014 JenneSummaryPSLPSL pointing monitoring

I'm not sure which pointing Rana wanted to fix today, but here's a measurement of the MC spots.  They actually look pretty good.  There is some room for improvement, but not a lot, so I'm leaving it alone for now, while I play with other things in the IFO.

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
[0.63368182839757914, 1.3004245778952557, 0.33621668795755993, -1.5585578137597658, -3.1344594013487286, -1.0533063060089816]

MCspots_6Jan2014.png

  9518   Fri Jan 3 18:21:45 2014 ranaSummaryPSLPSL pointing monitoring

 

 I went to the PSL table to re-align the input pointing to the IMC. After trying to optimize the pointing into the PMC and not succeeding I also then touched the wrong mirror and messed up our IOO QPD reference pointing.

The IMC is locking again, but I'll have to fix the pointing on Monday.

  9517   Fri Jan 3 15:19:39 2014 ranaSummaryPSLPSL pointing monitoring

 This is a 10-minute trend of the last 60 days of the pointing of the PSL beam.

The main fluctuation seems to be at the ~30 day time scale (not 24 hour) and its all in the vertical direction; the horizontal drift is ~10x less (as long as we believe there is no calibration error).

So what's causing all of this vertical shift? And why is there not just as much horizontal??

Attachment 1: PSL_pointing_2013.png
PSL_pointing_2013.png
  9516   Fri Jan 3 11:18:41 2014 SteveSummaryVACpower supply replaced with a short vent

Quote:

Quote:

 

 The temperature went down to room temp with temporary fan in the back. Voltage and current are stable.

Regardless, it will be replaced early next week.

Koji, Steve

 It was a bad experience again with our vacuum system.  The valves went crazy as we rebooted the computer. This was required for the swap in of a good 24V power supply.

The IFO was vented to 27 Torr through the annuloses, VA6, V7, Maglev,VM2 and VM1 (VC2 was open too)

I just opened the PSL shutter after a 4 hours pumpdown.

Condition: annuloses are not pumped, the IFO and the RGA are pumped as Atm2 shows

I will be here tomorrow morning to switch over to vacuum normal. 

More details later

 

 

 Events of the power supply swap:

1, Tested 24V DC ps from Todd

2, Closed V1, VM1 and all annulos valves to create safety net for the reboot. Turbo pumps left on running.

3, Turned computer off

4, Swap power supplies and turned it on

5, Turning the power on of c1vac2 created caos switching of valves. This resulted in a air vent as shown below.

6, VM1 was jammed and it was unable to close. The IOO beam shutter closed and the IFO was venting with air for a few minutes. Maglev did an emergency shut down. TP2's V4 and TP3' V5 closed. RP1 and RP3                           roughing         pumps turned on, their hose was not connected as usual. The RGA shut down to protect itself.

7, Closed annulos valves, stoped the vent at P1 27 torr as the vacuum control was  manually recovered

8, The Maglev and the annuloses were roughed out 500 mtorr . The Maglev was restarted.

9, The IFO pump down followed std procedure from 27 torr. VM1 was moving again as the pressure differential was removed from it.

 

 Remember: next time at atm .....rough down the cryo volume from 27 torr !

Attachment 1: rebootVENT.png
rebootVENT.png
  9515   Thu Jan 2 13:35:06 2014 KojiUpdateVAC vacuum monitor is still blank

We probably need to restart the machine, but I didn't want to touch c1vac1 and c1vac2. 

  9514   Thu Jan 2 10:50:24 2014 SteveUpdateVAC vacuum monitor is still blank

 The date is correct on this monitor.

Last stored RGA scan data Dec 20, 2013

IFO pressure at CC1 5.8e-6 Torr

Valve configuration: Vacuum Normal, confirmable only by manual checking of position indicators and pressure gauge controllers  readouts

 

Attachment 1: Help.png
Help.png
Attachment 2: vacation.png
vacation.png
Attachment 3: 20131220lastRGAscan.png
20131220lastRGAscan.png
  9513   Thu Jan 2 10:15:20 2014 JamieSummaryGenerallinux1 RAID crash & recovery

Well done Koji!  I'm very impressed with the sysadmin skillz.

  9512   Wed Jan 1 15:01:29 2014 KojiSummaryGeneralIFO recovery

IFO restart after the recovery of linux1

Machine recovery in the following order
- Start fb
- Start the following machines: mafalda, megatron, op340m
- Start c1ioo, c1lsc, c1sus, c1iscex, c1iscey

CDS recovery / burtrestore

- Confirm all of the RT systems are running in "green". If not, restart corresponding model.
- c1iscaux, ciscaux2 didn't have response (white boxes). Went to the LCS digital rack and power cycled these targets
- burtrestore: The snapshots at Dec 19 05:07 were used. For c1iscaux and c1iscaux2 the snapshots at Dec 22 05:07 were used.

fast machines
c1alsepics.snap
c1assepics.snap
c1asxepics.snap
c1calepics.snap
c1iooepics.snap
c1lscepics.snap
c1lspepics.snap
c1mcsepics.snap
c1oafepics.snap
c1pemepics.snap
c1rfmepics.snap
c1scxepics.snap
c1scyepics.snap
c1spxepics.snap
c1supepics.snap
c1susepics.snap
c1tstepics.snap

slow machines
c1auxex.snap
c1auxey.snap
c1aux.snap
c1iool0.snap
c1iscaux2.snap
c1iscaux.snap
c1psl.snap
c1susaux.snap

 

IFO recovery

- Reload watchdogs => restore sus damping
- MC misaligned but TEM00 was locked
- Gave a small touch on MC2 yaw => IMC almost aligned
- Autolocker wasn't running => Manually launched rather than wait for an hour for cron to launch it
- PMC was largely misaligned. => Aligned on the PSL table (PSLTRANS 0.640->0.753)
- MC WFS ON
- IFO X/Y arm locked and aligned with ASS.
- PRMI mode: manually aligned PRM. The PRMIsb momentally locked.

  9511   Tue Dec 31 23:19:58 2013 KojiSummaryGenerallinux1 RAID crash & recovery

Dec 22 between 6AM and 7AM, physical or logical failure has occure on the 4th disk in the RAID array on linux1.
This caused the RAID disk fell into the readonly mode. All of the hosts dependent on linux1 via NFS were affected by the incident.

Today the system has been recovered. The failed filesystem was restored by copying all of the files (1.3TB total) on the RAID to a 2TB SATA disk.
The depending hosts were restarted and we recovered elog/wiki access as well as the interferometer control system.


Recovery process

o Recover the access to linux1

- Connect an LCD display on the host. The keyboard is already connected and on the machine.
- One can login to linux1 from one of the virtual consoles, which can be switched by Alt+1/2/3 ...etc
- The device file of the RAID is /dev/sda1
- The boot didn't go straightforward as mounting of the disks accoding to /dev/fstab doesn't go well.
- The 40m root password was used to login with the filesystem recovery mode.
- Use the following command to make the editing of /etc/fstab available

# mount -o rw, remount /

- In order to make the normal reboot successfull, the line for the RAID in /etc/fstab needed to be commented out.

o Connect the external disk on linux1

- Brought a spare 2TB SATA disk from rossa.
- Connect the disk via an USB-SATA enclosure (dev/sdd1)
- Mount the 2TB disk on /tmpdisk
- Run the following command for the duplication

# rsync -aHuv --progress /home/ /tmpdisk/ >/rsync_KA_20131229_0230.log

- Because of the slow SCSI I/F, the copy rate was limited to ~6MB/s. The copy started on 27th and finished 31st.

o Restart linux1

- It was found that linux1 couldn't boot if the USB drive is connected.
- The machine has two SATA ports. These two are used for another RAID array that is not actually used. (/oldhome)
- linux1 was pulled out from the shelf in order to remove the two SATA disks.
- The 2TB disk was installed on the SATA port0.
- Restart linux1 but didn't start as the new disk is recognized as the boot disk.
- The BIOS setting was changed so that the 80GB PATA disk is recognized as the boot disk.
- The boot process fell into the filesystem recovery mode again. /etc/fstab was modified as follows.

/dev/VolGroup00/LogVol00 /                ext3    defaults        1 1
LABEL=/boot              /boot            ext3    defaults        1 2
devpts                   /dev/pts         devpts  gid=5,mode=620  0 0
tmpfs                    /dev/shm         tmpfs   defaults        0 0
proc                     /proc            proc    defaults        0 0
sysfs                    /sys             sysfs   defaults        0 0
/dev/VolGroup00/LogVol01 swap             swap    defaults        0 0
#/dev/md0                 /oldhome         ext3    defaults        0 1
/dev/sda1                /home            ext3    defaults        0 1
#/dev/sdb1                /tmpraid         ext3    defaults        0 1

- Another reboot make the operating system launched as usual.

o What's happen to the RAID?

- Hot removal of the disk #4.
- Hot plug of the disk #4.
- Disk #4 started to get rebuilt -> ~3hours rebuilding done
- This made the system marked as "clean". Now the raid (/dev/sdb1) can be mounted as usual.


o Nodus

- Root password of nodus is not known.
- Connect an LCD monitor and a Sun keyboard on nodus.
- Type Stop-A. This leads the nodus transition to the monitor mode.
- Type sync.
- This leads the system rebooted.

  9510   Sat Dec 21 10:53:35 2013 SteveUpdateVACpower supply replaced with a short vent-pumpdown completed

The recovery- pumpdown reached valve configuration  vacuum normal at 20 hours, cc1 7.7e-6 Torr

Lesson learned: turn all pumps off, close all valves before you reboot ! like you would prepare for AC power shut down.

 

Attachment 1: 20hrsVacNormal.png
20hrsVacNormal.png
  9509   Sat Dec 21 01:54:04 2013 KojiConfigurationLSCLSC Whitening for the DCPDs/CM servo replaced

The previous LSC whitening filters for the DCPDs were in an unknown state (although the transfer functions were actually measured and fit a while ago)
They had no DC gain control and some of the channels had modifications.

To make the setup clear, the filter module was replaced with the spare module without any modification.

The channels are now respoding to the whitening gain switches. So far there is no screen for the new  whitening gains yet.

Also I found that POX11 DC cable has not been connected. Now it is connected.

  9508   Fri Dec 20 23:00:41 2013 KojiUpdateVACpower supply replaced with a short vent

I'm leaving the 40m now. IFO is aligned. Everything look good.

- The main volume P1=5e-4, CC1=1.4e-5 is still pumped by TP1 and TP2

- RGA P4<0e-4, CC4 2.1e-7, is pumped by TP3

- The annuluses are isolated.

- RP1/2/3 are off.

  9507   Fri Dec 20 22:45:02 2013 KojiUpdateLSChigh bandwidth loop achieved for yarm

I checked the offset situation in the CM servo boost circuit. 

- The offset voltage on the CM servo screen is a raw DAC output. This number is diluted by the voltage divider before the amplifier.
  So, the actual offset of the boost circuit was much smaller. (~20mV)

- There is a offset trimmer on the board. This was adjusted so that the boost does not generate an output offset.

- So the default offset is 0V.

- When the arm was locked with (digital) POY11, the CM servo offset is necessary to be -2.7 (now).
  This means that analog POY11Q and digital POY11 has different offset for the best resonance transmission.
  That is believable if POY11I is contributing to the digital POY11 signal.

  9506   Fri Dec 20 20:04:01 2013 SteveUpdateVACpower supply replaced with a short vent

Quote:

Quote:

Instrument rack power supplies checked and labeled at present loads.

The vacuum rack Sorensen is running HOT! Their is only 0.3A load at 24V There is plenty of space around it.

It is alarming to me because all vacuum valve positions are controlled by this 24V

 The temperature went down to room temp with temporary fan in the back. Voltage and current are stable.

Regardless, it will be replaced early next week.

Koji, Steve

 It was a bad experience again with our vacuum system.  The valves went crazy as we rebooted the computer. This was required for the swap in of a good 24V power supply.

The IFO was vented to 27 Torr through the annuloses, VA6, V7, Maglev,VM2 and VM1 (VC2 was open too)

I just opened the PSL shutter after a 4 hours pumpdown.

Condition: annuloses are not pumped, the IFO and the RGA are pumped as Atm2 shows

I will be here tomorrow morning to switch over to vacuum normal. 

More details later

 

 

Attachment 1: 4hrPumpdown.png
4hrPumpdown.png
Attachment 2: pumpdownAfterHickup.png
pumpdownAfterHickup.png
Attachment 3: PSpumpdown.png
PSpumpdown.png
  9505   Fri Dec 20 18:00:02 2013 KojiSummaryCDSRCG parsing bug?

The bug is still there but the incorrect bits are now overridden.

Attachment 1: Screenshot-c1lsc-LSC.png
Screenshot-c1lsc-LSC.png
  9504   Fri Dec 20 17:41:25 2013 SteveUpdateGeneralProjector - lightbulb replaced

 The lamp lasted for 4,622 hours.

This time I purchased just the bare lamp itself . The housing doubles the price. The disadvantage of this technic that the lamp housing window can not be cleaned  perfectly. Atm2 picture is exaggerating this spot.

However, It does not degrade the image quality.

Roll over image to zoom in

      

 
 

Glamps RLC-061 Projector Original Bulb Lamp for VIEWSONIC Pro8200 Pro8300

 

 

Attachment 1: explodedlamp.jpg
explodedlamp.jpg
Attachment 2: clweanedWindowShield.jpg
clweanedWindowShield.jpg
  9503   Fri Dec 20 11:40:13 2013 JamieSummaryCDSRCG parsing bug?

I submitted a bug report for this:

https://bugzilla.ligo-wa.caltech.edu/bugzilla3/show_bug.cgi?id=553

However, given how old our RCG version is (2.5 vs. 2.8 current deployed at the sites) I don't think we're going to see any traction on this.  Even if this is still a bug in 2.8, they'll only fix it in 2.8.  There's no way they're going to make a bug fix release for 2.5.

We need to upgrade.

  9502   Fri Dec 20 10:08:43 2013 JamieConfigurationGeneralnetgpibdata is working again now

Quote:

Now netgpibdata is working again.

Usage:

cd /cvs/cds/rtcds/caltech/c1/scripts/general/netgpibdata   
./netgpibdata -i 192.168.113.108 -d AG4395A -a 10 -f meas01
./netgpibdata -i 192.168.113.105 -d SR785 -a 6 -f meas01   

Just wanted to point out that the correct "modern" path to this stuff is:

/opt/rtcds/caltech/c1/scripts/general/netgpibdata

This is, of course, the same directory, but under the correct "/opt/rtcds", instead of the old, incorrect "/cvs/cds".

  9501   Fri Dec 20 03:34:40 2013 KojiUpdateLSChigh bandwidth loop achieved for yarm

This too huge offset difference with/without "BOOST" switch should be checked.

  9500   Fri Dec 20 03:31:07 2013 KojiUpdateLSClock acquisition path for the CM servo

up/down scripts are to be made

(Offset Edit on Dec 20 10:38PM)


Configuration:
POY11QMon -> CM Servo In1 -> CM Servo -->Out1 -> ADC -> CM Slow FM -> LSC MC Servo FM -> ETMY(x1.0) -> DAC -> ETMY
                                       |
                                       -->Servo Out -> SR560 (DC, 1st order 30kHz LPF) -> MC In2


Lock acquisition path 1

Initial condition:

CM Slow FM:

  • Gain 2.6

CM Servo setting:

  • In1 Gain 31dB, SW:OFF, Offset -1.880, Boost Off, Super Boost Off, AO +8dB

MC Servo setting:

  • In2 10dB, SW:OFF

Acquisition:

  • Engage LSC
  • LSC MC servo gain +0.1, FM7/FM10 (Trigger FM2 with 3sec delay)
  • Turn on MC

Transition:

  • Enable AO path (CM servo In1 SW:ON, MC servo In2 SW:ON)
  • LSC MC gain +0.1 -> +0.2
  • AO path gain 8dB->14dB
  • LSC MC gain +0.2 -> +0.35
  • AO path gain 14dB->18dB
  • CM servo offset -1.88 -2.7 -> 0.8 0.0 (gradually)
  • Enable CM servo Boost

Lock acquisition path 2

Initial condition:

CM Slow FM:

  • Gain 2.6

CM Servo setting:

  • In1 Gain 31dB, SW:OFF, Offset -1.880, Boost Off, Super Boost Off, AO +20dB

MC Servo setting:

  • In2 10dB, SW:OFF

Acquisition:

  • Engage LSC
  • LSC Yarm G=+0.25 FM4/5 (Trigger FM3/6/7/8)

Transition:

  • Enable MC servo In2 (SW:ON)
  • Set LSC MC gain +0.2 FM7/10
  • Enable LSC MC (On)
  • Enable CM servo In1 (SW:ON)
  • Disable LSC Yarm (OFF)
  • Change CM servo offset -1.88 -> +0.700 -2.70 -> 0.0
  • Enable CM servo Boost
  • Turn on LSC CM FM2 (optional)

Transition to ETMY LSC to MCL

  • After all of the transition: LSC output matrix ETMY (+1.00)
  • LSC output matrix MC2 (-1.00)
  • LSC output matrix ETMY (0.00)
  9499   Fri Dec 20 01:24:11 2013 DenUpdateLSChigh bandwidth loop achieved for yarm

Koji, Den

CM Servo with POY11 successfully engaged. UGF: ~15kHz.


Tonight we decided to repeat one arm locking using high-bandwidth CM servo. We low-passed AO signal to avoid saturations of the EOM. We tried different configurations that compromise between noise and loop phase margin and ended up with a pole at 30kHz. SR560 is used as a low-pass filter.

Another problem that we faced was big (~2.6V) electronic offset at the input of 40:4000 BOOST. Once engaged, cavity would be kicked out of lock. We calibrated this offset to be almost half linewidth of the cavity (~300pm). To avoid lock loss during engaging the boost we increased common mode gain to maximum (31 dB).

Measured OL is attached. UGF is 15kHz, phase margin is 60 degrees. We have also simulated evolution of loop shape during bringing AO path. Plot is attached.

The final procedure is

  • set common gain up to 31dB, AO gain to 8dB, MC IN2 gain 10dB, CM offset 0.7V
  • lock arm with CM slow path with bandwidth of 200 Hz
  • enable AO path, gradually increase slow and fast gains by 12 dB
  • enable boost
Attachment 1: CM_OL_meas.pdf
CM_OL_meas.pdf
Attachment 2: cm_ol_sim.pdf
cm_ol_sim.pdf
Attachment 3: CM_slow_fast_cross.pdf
CM_slow_fast_cross.pdf
  9498   Fri Dec 20 00:16:39 2013 KojiSummaryCDSRCG parsing bug?

A while ago, I noticed that the most significant bits of the LSC whitening DOs are not toggling.
I track this issue down and found what is happening. I need experts' help.


To illuminate the issue, terminators are connected to Bit15 of the Bit2Word blocks in the LSC model (attached screen shots).

The corresponding source file is found in c1lsc.c at the following location.
The last channels of the Bit2Word are connected to lsc_cm_slow (the filter module).
This is the source of the issue. This wrong assignment of the connections
can't be changed by connecting Go-From tags to the chennels.

/opt/rtcds/caltech/c1/rtbuild/src/fe/c1lsc/c1lsc.c

3881// Bit2Word:  LSC_cdsBit2Word1                                                                                                              
3882{
3883double ins[16] = {
3884        lsc_as110_logicaloperator4,
3885        lsc_as110_logicaloperator1,
3886        lsc_refl11_logicaloperator4,
3887        lsc_refl11_logicaloperator1,
3888        lsc_pox11_logicaloperator4,
3889        lsc_pox11_logicaloperator1,
3890        lsc_poy11_logicaloperator4,
3891        lsc_poy11_logicaloperator1,
3892        lsc_refl33_logicaloperator4,
3893        lsc_refl33_logicaloperator1,
3894        lsc_pop22_logicaloperator4,
3895        lsc_pop22_logicaloperator1,
3896        lsc_pop110_logicaloperator4,
3897        lsc_pop110_logicaloperator1,
3898        lsc_as165_logicaloperator4,
3899        lsc_cm_slow
3900};
3901lsc_cdsbit2word1 = 0;
3902for (ii = 0; ii < 16; ii++)
3903{
3904if (ins[ii]) {
3905lsc_cdsbit2word1 += powers_of_2[ii];
3906}
3907}
3908}

3946// Bit2Word:  LSC_cdsBit2Word2                                                                                                              
3947{                                                                                                                                           
3948double ins[16] = {                                                                                                                          
3949        lsc_as55_logicaloperator4,                                                                                                          
3950        lsc_as55_logicaloperator1,                                                                                                          
3951        lsc_refl55_logicaloperator4,                                                                                                        
3952        lsc_refl55_logicaloperator1,                                                                                                        
3953        lsc_pop55_logicaloperator4,                                                                                                         
3954        lsc_pop55_logicaloperator1,                                                                                                         
3955        lsc_refl165_logicaloperator4,                                                                                                       
3956        lsc_refl165_logicaloperator1,                                                                                                       
3957        lsc_logicaloperator_cm_ctrl,                                                                                                        
3958        ground,                                                                                                                             
3959        ground,                                                                                                                             
3960        lsc_logicaloperator_popdc,                                                                                                          
3961        lsc_logicaloperator_poydc,                                                                                                          
3962        lsc_logicaloperator_poxdc,                                                                                                          
3963        lsc_logicaloperator_refldc,                                                                                                         
3964        lsc_cm_slow                                                                                                                         
3965};                                                                                                                                          
3966lsc_cdsbit2word2 = 0;                                                                                                                       
3967for (ii = 0; ii < 16; ii++)                                                                                                                 
3968{                                                                                                                                           
3969if (ins[ii]) {                                                                                                                              
3970lsc_cdsbit2word2 += powers_of_2[ii];
3971}
3972}
3973}

 

Attachment 1: Bit2Word1.png
Bit2Word1.png
Attachment 2: Bit2Word2.png
Bit2Word2.png
  9497   Thu Dec 19 21:16:16 2013 KojiConfigurationGeneralnetgpibdata is working again now

Now netgpibdata is working again.

Usage:

cd /cvs/cds/rtcds/caltech/c1/scripts/general/netgpibdata   
./netgpibdata -i 192.168.113.108 -d AG4395A -a 10 -f meas01
./netgpibdata -i 192.168.113.105 -d SR785 -a 6 -f meas01   

Jenne witnessed and certified that they were working fine.
Now no one can say "it does not work"!


These weeks I was annoyed by the fact that netgpibdata was messed up by dummies.
A terrible situation was found:

1. Someone pushed the factory reset of one of the wifi bridges (LINKSYS WET54G).
2. Someone gave wrong IPs to the bridges (192.168.1.* instead of 192.168.113.*)
3. Someone left a default IP to the bridges. This means the routers had the same IPs.

-------

I gave the IPs to the bridges. According lines of /etc/hosts in linux1 were updated.

192.168.113.230 WET54G1
192.168.113.231 WET54G2

All of the network settings are taped on the bridge now.
The reset switch of each bridge was covered by a tape so that dummies can't randomly push the button.

The command was tested with each device.

  9496   Thu Dec 19 19:45:12 2013 ericqUpdateGreen LockingX-Arm Green PDH Loop Stuff
With the fixed servo box, I remeasured the OLTF, the servo, and the low pass filter between the mixer output and servo input. Dividing the OLTF by the servo and LPF transfer functions should just leave the the [laser PZT->cavity->PD] transfer function, which should have the shape of the cavity pole plus any delay in the loop, up until the PZT is no longer linear / the measurement has bad SNR.

I'm missing a few pieces of the loop. While I know the PD gain in V/W, I don't know how much power is in the sideband, which affects the slope of the PDH error function. Also, I've found old ELOG posts mentioning either 1 or 5MHz/V being the NPRO PZT response, but am not sure how to determine what it actually is. These are essentially just scalars though, so finding the reason for low phase margin doesn't depend on them.

Here are the TFs I've measured ("residual" refers to OLTF/(servo*LPF)):



The teal "residual" TF presumably owes its shape to the cavity pole + the time delay around the loop. Messing around with the data, the shape fits very well to a real pole at 27kHz and a ~3usec delay. I have no real way to back that up as the unique truth behind it, however. Is there a good way to measure the delay? Without assuming any delay, the shape is best fit by a real pole at 26kHz and some funky complex zeros.

Another thing to look at is the CLG implied by the measurement of the OLTF, given by 1/(1-G). I plotted this quantity for the measured loop, and also for G/2 and 3G/2 to get an idea for how it changes as you turn the servo gain knob. I measured with the knob at 4.0. There seems to be quite a bit of gain peaking!



Also, I drew up a simple block diagram sort of thing to show how everything is connecting down at the green electronics rack at the end of the X arm (while totally glossing over the optical elements involved). This hopefully helps anyone who wants to inspect/take apart/massacre the setup.

  9495   Thu Dec 19 18:33:25 2013 GabrieleSummaryLSCEstimate of the sign of the PRC length error

Maybe I'm getting confused, but I still believe there is no way to decide the direction from yesterday's measurement.

Let's say for example that the arm sideband detuning from antiresonance is equivalent to a PRC length change of +1cm away from the position of ideal resonance of the sidebands without arms. Then we can get a measured separation of the sidebands, without arms, corresponding to 5cm both if the PRC is off by +4cm or by -6cm...

  9494   Thu Dec 19 14:40:42 2013 KojiUpdateCDSRFM Time over mitigation for c1mcs

I worked on the mitigation of c1mcs time-over issue this afternoon.

The timing for the c1mcs is successfully reduced from >60us to 45us.


The previous models are svned in redoubt as follows:

MCS rev. 6696
RFM rev. 6697
IOO rev. 6698

What I changed was:

- Remove connection from ALS (on c1ioo) to MCS (on c1sus). This should be all done in LSC. (# of RFM IPC in MCS -1)

- MC2 trans QPD filters are moved from IOO to MCS to reduce the RFM channels in MCS.
  Previously the signals for the 4 segments are sent. Now the processed siganls (pit/yaw/sum) are sent. (# of RFM IPC in IOO -1, MCS -1)

- WFS MC3 feedback channels are moved from MCS to RFM to distribute the RFM channels (# of RFM IPC in MCS -2, in RFM +2)

model    prev. timing[us] current timing[us]  diff in time[us]  diff in ch#
c1mcs         >60                45                -15              -4
c1rfm         47                 53                + 6              +2       
c1ioo         47                 36                -11              -1

Revisions of the new models:
MCS rev. 6702
RFM rev. 6701
IOO rev. 6700

  9493   Thu Dec 19 12:51:57 2013 JenneSummaryLSCEstimate of the sign of the PRC length error

My hunch is that the PRC is SHORT by a few cm, not long. 

In my Optickle simulation, the sidebands are not perfectly co-resonating in the PRMI when the arms are not locked.  See Fig 1, which is the fields in the PRMI using the design PRC length.  If I add 5cm to the PRC length, I get Fig 2, where the peaks are about the same separation, but the upper and lower sidebands have swapped sides of the 0 mark.  However, if I remove 5cm from the PRC length, I get Fig 3, where the peaks are much farther apart than in Fig 1.  This case looks more similar to the data that Gabriele plotted in his elog entry, where the peaks are separated by at least a linewidth.  This is not at all conclusive, but it's a guess for which direction we need to move.  Obviously doing an actual measurement will be better. 

My tummy feelings also agree with this simulation:  When we flipped PR3 (the only optic change in the PRC since Gabriele and I measured the 55MHz peak separation in April), since the HR side of the optic is now at the back, we had to push the whole suspension cage forward to get the beam aligned to the Yarm.  Conversely, however, transmitting through the glass substrate adds to the optical path length.  So, my tummy feelings may be wrong.

Figure 1, PRC at design length, PRMI sweep:

DesignLength.png

Figure 2, PRC 5cm longer than design length:

Plus5cm.png

Figure 3, PRC 5cm shorter than design length:

Minus5cm.png

 

  9492   Thu Dec 19 03:29:34 2013 DenUpdateLSCCM servo test using yarm is complete

Koji, Den

Procedure:

  • lock yarm on IR, wire POY to CM input
  • transition arm to CM length path by actuating on IMC
  • increase AO gain for a stable crossover
  • engage CM boosts

Result:

  • arm can be kept on resonance and even acquired on MC2
  • stable length / AO crossover is achieved
  • high bandwidth loop can not be engaged because POY signal is too noisy and EOM is running out of range

We spent some time tuning CM slow servo such that fast path would be stable in the AO gain range -32db -> 29dB (UGF=20kHz) when all boosts are turned off and common gain is 25dB. Current filters that we use for locking are not good enough - AO can not be engaged due to oscillations around 1kHz. This is clearly seen from slow path closed loop transfer function. I will attach servo shapes tomorrow.

Attached plot "EOM" shows EOM rms voltage while changing AO gain from -10dB to 4dB. For UGF of 20kHz we need AO gain of 29dB.

It seems we can start using CM servo for CARM offset but the sensor should be at least factor of 30 better than POY. Add another factor of 10 if we would like to use BOOST 2 and BOOST 3.

Attachment 1: EOM.png
EOM.png
  9491   Wed Dec 18 18:45:39 2013 JenneUpdateLSCREFL 165 demod phases

I checked out the REFL165 demod phase, and it looks like it was okay.  it was -20.9 degrees.  I turned on my sensing matrix oscillators, and maximized the PRCL signal in the REFL165_I_ERR channel, and got a pretty good maximization at +155 degrees.  I used this to lock the PRMI on sidebands, with MICH gain of +0.3, and PRCL gain of +0.1 . 

Since this is working, I'm leaving the REFL 165 phase here, at +155 degrees, although this is almost exactly 180 degrees from what Den left it at, so I'm not sure why I was not able to lock with a demod phase of -20.9.  (I tried all 4 permutations of signs, with gain values of the same magnitude (0.3 for MICH, 0.1 for PRCL), and wasn't able to lock.  I'll try to figure this out tomorrow, but it was time for the meeting, then the IFO has been busy doing more important things the rest of the afternoon.

Plan for checking:  Lock with demod phase of 155, measure TF to one of the other REFL diodes (11, 33 or 55), lock on that other REFL diode.  Then, change the REFL 165 demod phase back to -20.9, and measure the transfer function again.  Hopefully the answer is just that I was doing something dumb, and it works easily.  This test/measurement should only take a few minutes, but it'll make me happier knowing that things still work as they should.

  9490   Wed Dec 18 17:28:50 2013 GabrieleSummaryLSCEstimate of the PRC length error

Measurement 

Looking back at what I did in april (see log #8411) I realized that it is possible to get an estimate of how much the PRC length is wrong looking at the splitting of the sideband resonant peak as visible in the POP_110_I signal. With the help of Jenne the PRMI was aligned and left swinging. The first plot shows a typical example of the peak splitting of 55MHz sidebands. This is much larger than what was observed in April.

When the sidebands resonate inside the PRC they get a differential dephasing given by 

dPhi = 4*pi*f_mod/c * dL

where dL is the cavity length error with respect to the one that makes the sidebands perfectly resonant when the arms are not there. This is not exactly the error we are interested in, since we should take into account the shift from anti-resonance of the SBs in the arm cavities.

Nevertheless, I can measure the splitting of the peak in units of the peaks full width at half maximum (FWHM). I did this fitting few peaks with the sum of two Airy peaks. Here is an example of the result

data_fit2.png

The average splitting is 1.8 times the FWHM. Knowing the PRC finesse, one can determine the length error:

dL = c / (4 * f_mod * Finesse) * (dPhi / FWHM)

Assuming a finesse of 60, I get a length error of 4 cm.

To get another estimate, we kicked the PRM in order to get some almost linear sweeps of the PRC length. Here is one of the best results:

data_kick.png

The distance between consecutive peaks is the free spectral range (FSR) of the PRC cavity. Again, I can measure the peak splitting in units of the FSR and determine the length error:

dL = c / (4 * f_mod) * (dPhi / FSR)

The result is again a length error of 4 cm.

Simulation

An error of 4 cm seems pretty big. Therefore I set up a quick simulation with MIST to check if this makes sense. Indeed, if I simulate a PRMI with the 40m parameters and move the PRC length from the optimal one, I get the following result for POP_110_I, which is consistent with the measurement.

pop110_vs_dL.png

 

 Therefore, we can quite confidently assume that the PRC is off by 4 cm with respect to the position that would make the 55 MHz sideband resonant without arms. Unfortunately, it is not possible with this technique to infere the direction of the error.

 

 

 

 

 

 

Attachment 3: pop110_vs_dL.png
pop110_vs_dL.png
  9489   Wed Dec 18 15:35:48 2013 SteveUpdatePEMexcess seismic noise

This is not a test. Life can be dangerous in the 40m Control  Room.

Attachment 1: seismicCar.jpg
seismicCar.jpg
  9488   Wed Dec 18 13:34:03 2013 SteveUpdateGeneralLIGOX people

40m crew and visitor Holger Muller from Berkeley.

Attachment 1: 40m2013Dec.jpg
40m2013Dec.jpg
Attachment 2: 40mCup.jpg
40mCup.jpg
  9487   Wed Dec 18 11:37:12 2013 JenneUpdateLSCPOP22 and POP110 demod phases

Somehow the POP22 and POP110 demod phases weren't correct anymore.  I guess Den saw this after he changed the setup for the REFL165 PD at the LSC rack, but didn't elog it. 

I went out to the LSC rack, and found that the power supply that is supplying the amplifiers for both POP22/110 and REFL165 was set to ~16V each channel.  I put it back to 15V for each channel.  I don't know what Den intended for the 165 amplifier (more volts is more gain), but the POP22/110 amplifier usually runs with 15V. 

I also reset the POP22 and POP110 demod phases.  Since I'm not able to lock PRMI on sideband this morning (why?!?!), I locked on the carrier, and moved the phases around until POP22 and POP110 were both maximally negative.  The phases are/were:

  OLD [deg] NEW [deg]
POP22 107.2 -165.0
POP110 95.0 150.0

This is a ~60 degree change for both PDs. 

I am not sure if Den ever checked the demod phase of REFL165 after he put in the new SMA cable (there's no mention of it in the elog!), so I'm going to check that to see if it helps get PRMI locking back.  I know that Den had also been using REFL11 for PRMI locking, but the parameters he used for that aren't in the log either.

  9486   Wed Dec 18 11:32:34 2013 ranaUpdateLSCXend QPD schematic investigation

 

 Since we use the TransMon QPD for triggering the high/low gain switching we need to run with the whitening OFF during lock acquisition and the turn it on after we have the arms locked with ALS. This should be put into the up/down scripts.

  9485   Wed Dec 18 03:29:48 2013 DenUpdateLSCyarm locked on mc

As a CM slow path test I locked free swinging yarm by actuating on MC length with bandwidth of 200 Hz. Crossover with AO is not stable so far.

I used xarm as an ool frequency noise sensor. MC2 violin mode is at 645 Hz, I have added a notch filter to LSC-MC2 bank.

Attachment 1: MC_ARM.pdf
MC_ARM.pdf
  9484   Wed Dec 18 00:26:15 2013 JenneUpdateLSCXend QPD schematic investigation

I have looked at the photo of the Xend QPD from elog 9367, as well as the schematic for the board (D990272). 

Things that will need swapping out:

  • Thick film resistors in the signal path need to be changed to thin film.
  • MAX333 needs to be replaced with MAX333A.  The 333 has "ON Resistance" of 140-175 ohms, whereas the 333A has "ON Resistance" of 20-35 ohms.
  • AD797 needs to be replaced by OPA140.  The 797 is a low voltage noise op-amp, but for a diode we want low current noise.  AD797 has 2pA/rtHz at 1kHz, whereas the OP140 has 0.8fA/rtHz at 1kHz (see Zach's elog 8125 re: OPA140).

I have ordered from digikey via techmart 10 each of the MAX333A's and the OPA140's.  (4 per QPD times 2 QPDs plus 2 spares = 10).  Both of these new chips have the same footprint and pinout as the part that they are replacing, so it'll be a fairly easy task.

Next up, I need to make a LISO model for the circuit for one of the quadrants, to see what shape it'll turn out to be.  Part of this will include deciding what resistors and capacitors to put in the OPA140 gain stage. 

Right now, the AD797s say on the schematic that the gain options are different by a factor of 5, but the actual QPD has a different resistor than is on the schematic, and there is also a capacitor in parallel with each resistor, so I need to just pull those out, and pick some values that make sense to me. 

Rana and I have discussed ignoring the 2nd and 3rd gain switching options on each quadrant, as that is getting to be more fine control than we need. 

Other things on the board:

  • The 50 ohm resistors to ground for the "QPD_rtn" have all shorted.  Rana says this is good, so leave it as-is.
  • The positive input to the AD797's all have a 100 ohm resistor to ground, rather than just being connected to ground.  Why is this??

For now, I will probably just work on the QPD head, and not the whitening board.  For now, we can run with 1 stage of whitening, and if we need lower noise, we can revisit the whitening board (including replacing the thick film resistors with thin film).

When thinking about what gains I want on my gain stages, I want to have my full arm power (~700 TRX units) be ~20,000 counts from the ADC.  So, I want my single arm power (1 TRX unit) to be ~30 counts from the ADC.  This is not such a big number, so this may also require more thinking. 

ELOG V3.1.3-